Circular No. 8511

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005bi AND 2005bj T. Boles, Coddenham, England, reports the discovery of two apparent

supernovae on unfiltered CCD images: SN 2005bi (mag 18.5) was observed on Apr. 11.015 and 12.004 UT at α = 16^{h}27^{m}21^{s}.27, δ = +39^{o}06′43′′.2 (equinox 2000.0), which is ≈ 10′′.0 west and 8′′.0 north of the center of MCG +07-34-36; nothing was visible at this location on his images taken on 2004 Aug. 6 and May 27 (limiting mag 19.5) or on Digitized Sky Survey plates from 1993 (limiting red mag 20.5) and 1989 (limiting blue mag 20.5). SN 2005bj (mag 17.7) was observed on Apr. 11.096 and 12.014 at α = 16^{h}49^{m}44^{s}.74, δ = +17^{o}51′48′′.7, which is ≈ 12′′.0 east and 3′′.5 south of the center of MCG +03-43-5; nothing was visible at this location on his images taken on 2004 July 5 and May 22 (limiting mag 19.5) or on DSS plates from 1992 (limiting red mag 21.0) and 1991 (limiting blue mag 20.5).

(5905) JOHNSON B. Warner, Colorado Springs, CO; P. Pravec and P. Kušnirák, Ondře-

jov Observatory; D. Pray, Coventry, RI; A. Galád and Š. Gajdoš, Modra Observatory; and P. Brown and Z. Krzeminski, Department of Physics and Astronomy, University of Western Ontario, report that photometric observations obtained during 2005 Apr. 1–11 show that (5905) is a binary system with an orbital period of 21.78 hr. The primary rotates with a period of 3.783 hr, and its lightcurve amplitude of 0.11 mag is indicative of a nearly spheroidal shape. Mutual events 0.15–0.18 mag deep indicate a secondary-to-primary mean-diameter ratio of 0.4.

V2361 CYGNI M. Bode, R. Zamanov, and J. Marchant, Liverpool John Moores Uni-

versity; and T. J. O’Brien, Jodrell Bank Observatory, report the following CCD magnitude estimates of this nova, obtained with the 2-m Liverpool telescope on La Palma: Mar. 17.229 UT, V = 18.80, r′ = 17.25, i′ = 16.64; 18.210, 18.80, 17.25, 16.65; 29.181, 18.55, 16.77, 16.51; 30.179, 18.30, 16.72, 16.40; Apr. 2.170, 18.38, 16.67, 16.53. Uncertainties due to nearby star are ∼ 0.10 mag in Johnson V, and 0.05 mag in Sloan r′ and i′.

SUPERNOVA 2005be T. Puckett writes that he has revised the unfiltered magnitudes of SN

2005be reported on IAUC 8506 as follows: Apr. 5.27 UT, 16.1; 6.21, 16.4.

2005 April 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green